# WIMP/FIMP dark matter and primordial black holes with memory burden effect

### v1

@@ 3.1 WIMPs and PBHs
First, we consider DM made of WIMPs and PBHs. FO of WIMPs occurs at [2]
@@
where $T_{\rm FO}$ is the FO temperature. We assume that FO occurs during a radiation-dominated era. The Hubble parameter is related to the radiation temperature and time by relation $H=(\pi^{2}g_{\ast}/30)T^{2}/(3M_{\rm P})=1/(2t)$. Thus, the FO time is obtained as follows:
@@
The relic abundance of WIMPs $\chi$ via FO is calculated as [38]
@@
where $\langle\sigma v\rangle$ is the thermally averaged cross-section for $\chi\chi\rightarrow\bar{f}f$ and $s_{0}=2.89\times 10^{3}\ {\rm cm^{-3}}$ and $\rho_{\rm c}=3H_{0}^{2}/(8\pi G)=1.05h^{2}\times 10^{-5}\ {\rm GeV\ cm^{-3}}$ are the entropy and critical densities in the present day, respectively. Considering the so-called WIMP miracle, we parameterize
@@
where $\alpha_{\rm WIMP}^{2}$ is the effective coupling for $\chi\chi\leftrightarrow\bar{f}f$. The relic abundance of WIMPs is given by
@@
If the FO mechanism alone allows to produce DM, observed relic abundance $\Omega_{\rm DM}h^{2}=0.12$ [49] is consistent with $\alpha_{\rm WIMP}\simeq 0.026$ and $m_{\chi}\simeq 1$ TeV.
If the PBH emission of WIMP $\chi$ precedes FO of $\chi$, the WIMPs emitted by Hawking radiation may reach chemical equilibrium with bath particles due to the massive thermal production of WIMPs. In this case, WIMPs from PBH evaporation provide no additional contribution to the relic abundance of thermally produced WIMPs. In contrast, if PBHs evaporate after WIMP FO, WIMPs emitted from PBHs may neither thermalize with the bath particles nor annihilate with each other. In this case, the WIMPs produced from PBHs may contribute to the relic abundance of DM [38, 40].
@@
We assume that the DM particle is a Majorana fermion (e.g., $\xi=3/4$ and $g_{\chi}=1$). In addition, we omit the negligible contribution of $g_{\chi}=1$ to the effective number of degrees of freedom and use $g_{*}=106.75$. The present-day density of DM can be obtained as follows:
$\Omega_{\rm DM}h^{2}$ =$=\begin{cases}\Omega_{\rm FO}h^{2}+\Omega_{\rm PBH}h^{2}&{\rm for}\ t_{\rm ev}^{k}<t_{\rm FO}\\ \Omega_{\rm FO}h^{2}+\Omega_{\rm PBH}h^{2}+\Omega_{\rm ev}h^{2}&{\rm for}\ t_{\rm FO}<t_{\rm ev}^{k}\\ \end{cases},$
$=1.7\times 10^{-10}\frac{1}{\alpha_{\rm WIMP}^{2}}\left(\frac{m_{\chi}}{\rm GeV}\right)^{2}+4.2\times 10^{26}q\beta\left(\frac{M_{\rm P}}{M_{\rm in}}\right)^{1/2}$
+$\quad+\begin{cases}0&{\rm for}\ t_{\rm ev}^{k}<t_{\rm FO}\\ 2.3\times 10^{5}C\beta\frac{m_{\chi}}{\rm GeV}&{\rm for}\ t_{\rm FO}<t_{\rm ev}^{k}\\ \end{cases},$ (31)
@@
where the first–third terms are due to FO, survived PBHs, and PBH evaporation, respectively. Coefficient $C$ is defined as
$C=\begin{cases}(1-q^{2})\left(\frac{M_{\rm in}}{M_{\rm P}}\right)^{1/2},&{\rm for}\ m_{\chi}<T_{\rm in}\\ \left(\frac{M_{\rm P}}{M_{\rm in}}\right)^{3/2}\left[\left(\frac{M_{\rm P}}{m_{\chi}}\right)^{2}-\left(\frac{qM_{\rm in}}{M_{\rm P}}\right)^{2}\right],&{\rm for}\ m_{\chi}>T_{\rm in}\\ \end{cases}$ (32)
@@
We study the case of $\Omega_{\rm FO}h^{2}\gg\Omega_{\rm ev}h^{2}$. In the following numerical calculations, we assume $k=1$ as the simplest case. As a benchmark, we consider $\alpha_{\rm WIMP}=0.0323$ and $m_{\chi}=700$ GeV. In this case, we obtain $\Omega_{\rm FO}h^{2}=0.08$. The left panel in Fig. 1 shows the relic abundance of WIMP due to Hawking radiation from PBHs $\Omega_{\rm ev}h^{2}$ according to the initial PBH mass, $M_{\rm in}$, for initial PBH density $\beta=1\times 10^{-20}$. The blue solid curve indicates prediction without the memory burden effect. The dotted curves show the prediction with the memory burden effect for various $q$ values. Relation $\Omega_{\rm FO}h^{2}\gg\Omega_{\rm ev}h^{2}$ is satisfied for $q=0.5-5\times 10^{-6}$. As the relic density of WIMPs via Hawking radiation should vanish for $t_{\rm ev}^{k}<t_{\rm FO}$, a lower limit of $M_{\rm in}$ appears in each curve. These lower limits can indicate shifts to a lighter mass owing to the memory burden effect via relation $t_{\rm ev}^{k}\simeq 1/\Gamma_{\rm BH}^{k}\propto(qM_{\rm in})^{3+2k}$ from Eqs. (12) and (13).
@@
Figure 1: Relic abundance of DM considering WIMPs and PBHs. Relic abundance of WIMPs due to Hawking radiation from PBHs $\Omega_{\rm ev}h^{2}$ (left) and total relic abundance of DM $\Omega_{\rm DM}h^{2}$ (right) according to initial PBH mass $M_{\rm in}$.
Figure 2: Constraints on initial PBH density $\beta$ considering WIMPs and PBHs. Top left: $\beta$ and $\beta_{\rm c}$ according to $M_{\rm in}$ for $q=5\times 10^{-6}$. Top right: $\beta$ according to $M_{\rm in}$ for various $q$ values. Bottom left: $\beta$ according to mass of WIMPs, $m_{\chi}$. Bottom right: $\beta$ according to coupling of WIMPs, $\alpha_{\rm WIMP}$.
@@
The right panel in Fig. 1 shows the total relic abundance of DM, $\Omega_{\rm DM}h^{2}$, according to the initial PBH mass, $M_{\rm in}$. As for the left panel, the blue solid curve indicates prediction without the memory burden effect. The dotted black horizontal line indicates the observed relic abundance, $\Omega_{\rm DM}h^{2}=0.12$. The remaining dotted curves show the prediction with the memory burden effect for various $q$ values. The relic density of WIMPs via FO is fixed at $\Omega_{\rm FO}h^{2}=0.08$, and relation $\Omega_{\rm DM}h^{2}\simeq\Omega_{\rm FO}h^{2}+\Omega_{\rm PBH}h^{2}$ is satisfied. Hence, the trends in the dotted curves are reflected by the relic density of survived PBHs with the memory burden effect, $\Omega_{\rm PBH}h^{2}\propto q\beta/M_{\rm in}^{1/2}$. The upper limit of $M_{\rm in}$ per dotted curve is due to requirement $\beta<\beta_{\rm c}\propto 1/(q^{5/2+k}M_{\rm in}^{1+k})$ (see also the top-left panel in Fig. 2). The relic abundance of DM is highly sensitive to the memory burden effect. For example, a small memory burden effect of $q\lesssim\mathcal{O}(10^{-3})$ is only allowed for $M_{\rm in}\lesssim\mathcal{O}(10^{6})$ g, $\beta=1\times 10^{-20}$, and $\Omega_{\rm FO}h^{2}=0.08$.
@@
Figure 2 show the constraint on the initial PBH density, $\beta$. The top-left panel shows $\beta$ and $\beta_{\rm c}$ according to $M_{\rm in}$ for $q=5\times 10^{-6}$ as an example. This panel shows the upper limit of $M_{\rm in}$ with $\beta<\beta_{\rm c}$ for $q=5\times 10^{-6}$. The top-right panel shows $\beta$ according to $M_{\rm in}$ for various $q$ values. The upper limit of $M_{\rm in}$ in each curve is due to requirement $\beta<\beta_{\rm c}$. The bottom-left panel shows $\beta$ according to the mass of WIMPs, $m_{\chi}$. The upper limit of mass $m_{\chi}$ is obtained at $\Omega_{\rm DM}h^{2}=\Omega_{\rm FO}h^{2}$. As we fix $\Omega_{\rm DM}h^{2}=0.12$ and $\alpha_{\rm WIMP}=0.0323$, the upper limit, $m_{\chi}=858$ GeV, is obtained using Eq. (30). The bottom-right panel shows $\beta$ according to the coupling of WIMPs, $\alpha_{\rm WIMP}$. Like for $\beta$ according to $m_{\chi}$, the lower limit of $\alpha_{\rm WIMP}$ is caused by the relation in Eq. (30). As a strong memory burden effect yields long-lived PBHs, the initial PBH density, $\beta$, for $\Omega_{\rm DM}h^{2}=0.12$ decreases with increasing $q$ in all panels, except for the top-left panel in Fig. 2.
@@ 3.2 FIMPs and PBHs
Figure 3: Relic abundance of DM and constraint on initial PBH density for FIMPs and PBHs. Only figures that differ from the case considering WIMPs are shown. Top: relic abundance of FIMPs owing to Hawking radiation from PBHs, $\Omega_{\rm ev}h^{2}$, according to initial PBH mass $M_{\rm in}$. Bottom left: $\beta$ according to mass of FIMPs, $m_{\chi}$. Bottom right: $\beta$ according to coupling of FIMPs, $\alpha_{\rm FIMP}$.
Next, we consider DM made of FIMPs and PBHs. We assume that the masses of bath particles are negligible compared with those of FIMPs and that the FI mechanism is described by a pair production process with constant matrix element $|\mathcal{M}(f_{1}f_{2}\rightarrow\chi\chi)|^{2}=\alpha_{\rm FIMP}^{2}$. For a simple model of FIMPs [50, 38], the relic abundance of FIMPs, $\chi$, via FI is given by [50, 38]
@@
If the FI mechanism alone allows to produce DM, the observed relic abundance of DM, $\Omega_{\rm DM}h^{2}=0.12$, is consistent with $\alpha_{\rm FIMP}=3.52\times 10^{-11}$ [50].
Given the feeble interactions of FIMPs, the FIMPs emitted by Hawking radiation can always contribute to the relic abundance of DM. The present-day DM density is given by
@@
$=9.7\times 10^{19}\alpha_{\rm FIMP}^{2}+4.2\times 10^{26}q\beta\left(\frac{M_{\rm P}}{M_{\rm in}}\right)^{1/2}+2.3\times 10^{5}C\beta\frac{m_{\chi}}{\rm GeV},$ (34)
where the first–third terms are due to FI, survived PBHs, and PBH evaporation, respectively.
Like for the WIMP case, we consider $\Omega_{\rm FI}h^{2}\gg\Omega_{\rm ev}h^{2}$. As a benchmark, we consider $\alpha_{\rm FIIMP}=2.872\times 10^{-11}$, $m_{\chi}=700$ GeV ($\Omega_{\rm FI}h^{2}=0.08$), and $\beta=1\times 10^{-20}$. For this benchmark, we obtain the same prediction of $\Omega_{\rm DM}h^{2}$ with $M_{\rm in}$, as shown in the right panel of Fig. 1. In addition, the prediction of $\beta$ with $M_{\rm in}$ for this benchmark considering FIMPs becomes the same as that considering WIMPs in Fig. 2. Thus, we omit the plots of $\Omega_{\rm DM}h^{2}$ $M_{\rm in}$ and $\beta$ according to $M_{\rm in}$ for the FIMP case.
The top panel in Fig. 3 shows the relic abundance of FIMPs due to Hawking radiation from PBHs, $\Omega_{\rm ev}h^{2}$, according to initial PBH mass $M_{\rm in}$ for $\beta=1\times 10^{-20}$. $\Omega_{\rm FO}h^{2}\gg\Omega_{\rm ev}h^{2}$ is satisfied for $q=0.5-5\times 10^{-6}$. Unlike the WIMP case, $\Omega_{\rm ev}h^{2}$ can always contribute to the relic abundance with no lower limit of $M_{\rm in}$.
The two bottom panels in Fig. 3 show the constraints on initial PBH density $\beta$ considering FIMPs. The bottom-left panel shows $\beta$ according to the mass of FIMPs, $m_{\chi}$. Unlike the WIMP case, no strong constraint on the mass of FIMPs is observed at $\Omega_{\rm DM}h^{2}=\Omega_{\rm FO}h^{2}$. The bottom-right panel shows $\beta$ according to the coupling of FIMPs, $\alpha_{\rm FIMP}$. Unlike the WIMP case, the upper limit of $\alpha_{\rm FIMP}$ is caused by the relation in Eq. (33).
Like for the WIMP case, a strong memory burden effect yields long-lived PBHs, and initial PBH density $\beta$ for $\Omega_{\rm DM}h^{2}=0.12$ decreases with increasing $q$ in the two bottom panels of Fig. 3.

### v3

@@ 4.1 WIMPs and PBHs
First, we consider DM made of WIMPs and PBHs. The FO of WIMPs occurs at [2]
@@
where $T_{\rm FO}$ is the FO temperature. We assume that FO occurs during a radiation-dominated era. The Hubble parameter is related to the radiation temperature and time as $H=(\pi^{2}g_{\ast}/30)T^{2}/(3M_{\rm P})=1/(2t)$. Thus, the FO time is
@@
The relic abundance of WIMPs due to FO, $\chi$, is [48]
@@
where $\langle\sigma v\rangle$ is the thermally averaged cross section at $\chi\chi\rightarrow\bar{f}f$ and $s_{0}=2.89\times 10^{3}\ {\rm cm^{-3}}$ is the entropy density. Considering the so-called WIMP miracle, we parameterize
@@
where $\alpha_{\rm WIMP}^{2}$ is the effective coupling at $\chi\chi\leftrightarrow\bar{f}f$. The relic abundance of WIMPs is
@@
If the FO mechanism alone produces DM, then the observed relic abundance, $\Omega_{\rm DM}h^{2}=0.12$ [110], is consistent with $\alpha_{\rm WIMP}\simeq 0.026$ and $m_{\chi}\simeq 1$ TeV.
The potential impact of PBH evaporation without the memory burden effect on the production of DM particles via thermal processes has been deeply investigated. In the context of the FO or FI mechanism, the authors in Ref. [50] considered fermionic DM interacting with SM particles via a spin-1 mediator. They focused on a case where the mediator is heavier than the DM but has a smaller mass than the reheating temperature.
In the case of the FO mechanism, the authors confirmed that DM particles produced from PBH evaporation thermalize with SM particles if the PBHs evaporate before FO. Therefore, PBHs do not affect the FO mechanism dynamics or contribute to the final relic abundance. Conversely, when PBH evaporation occurs after FO, two scenarios are possible.
In the first scenario, PBHs may dominate the universe’s energy density before evaporating. With PBHs dominating the universe’s energy density, PBHs act as a DM source by continuously injecting entropy into the visible sector, possibly affecting thermal production. If FO occurs during or before a PBH-dominated era, then PBH evaporation can alter standard FO predictions during PBH domination. However, unless particles from PBHs interact efficiently with thermal DM remnants and destroy FO predictions, PBHs cannot simultaneously affect FO production dynamics and contribute significantly to the relic density if they evaporate after FO. This is due to the presence of the mediator, which provides additional annihilation channels for DM and causes resonant interactions between DM and SM particles. The interaction between the evaporated DM particles and the preexisting relic abundance or between the evaporated DM particles and SM particles must be efficient for PBHs to dominate the universe’s energy density before evaporating. In this case, the Boltzmann equation should be solved at the level of phase-space distributions because of the nontrivial phase-space distribution of the evaporation products [50].
In the second scenario, PBHs never dominate the energy density. In this case, the FO mechanism and PBH evaporation can both contribute without affecting each other.
In this study, we focus on a case where the effect of the thermal production of WIMPs/FIMPs considerably exceeds that of PBH particle production via Hawking radiation. In addition, we assume that no early-matter-dominated era exists and PBHs never dominate the energy density in the history of the universe (we focus on a scenario where $\beta<\beta_{\rm c}$ holds). Therefore, DM particles from PBH evaporation are expected not to interact with thermally produced WIMPs.
To ensure that thermalization does not occur, we evaluate the interaction rates between particles from PBHs and thermally produced particles and compare them with the Hubble scale, as done in Ref. [50]. However, the current study does not examine the details of this problem. We take the range of $q$ at $\Omega_{\rm FO}h^{2}\gg\Omega_{\rm ev}h^{2}$.
@@
We assume that the DM particle is a Majorana fermion (e.g., $\xi=3/4$ and $g_{\chi}=1$). In addition, we omit the negligible contribution of $g_{\chi}=1$ to the effective number of degrees of freedom and use $g_{*}=106.75$. The present-day density of DM is
$\Omega_{\rm DM}h^{2}$ $=\begin{cases}\Omega_{\rm FO}h^{2}+\Omega_{\rm PBH}h^{2}&{\rm for}\ t_{\rm ev}^{k}<t_{\rm FO}\\ \Omega_{\rm FO}h^{2}+\Omega_{\rm PBH}h^{2}+\Omega_{\rm ev}h^{2}&{\rm for}\ t_{\rm FO}<t_{\rm ev}^{k}\\ \end{cases},$
$=1.7\times 10^{-10}\frac{1}{\alpha_{\rm WIMP}^{2}}\left(\frac{m_{\chi}}{\rm GeV}\right)^{2}$
$\quad+4.2\times 10^{26}q\beta\left(\frac{M_{\rm P}}{M_{\rm in}}\right)^{1/2}$
$\quad+\begin{cases}0&{\rm for}\ t_{\rm ev}^{k}<t_{\rm FO}\\ 2.3\times 10^{5}C\beta\frac{m_{\chi}}{\rm GeV}&{\rm for}\ t_{\rm FO}<t_{\rm ev}^{k}\\ \end{cases},$ (39)
@@
where the first–third terms are due to FO, survived PBHs, and PBH evaporation. The coefficient $C$ is
[…]
@@
We study the case of $\Omega_{\rm FO}h^{2}\gg\Omega_{\rm ev}h^{2}$. In the following numerical calculations, we assume $k=1$ as the simplest case. We regard $\alpha_{\rm WIMP}=0.0323$ and $m_{\chi}=700$ GeV as the benchmark. In this case, we obtain $\Omega_{\rm FO}h^{2}=0.08$.
@@
Figure 2: Relic abundance of DM considering WIMPs and PBHs. Left: relic abundance of WIMPs due to Hawking radiation from PBHs, $\Omega_{\rm ev}h^{2}$. Right: total relic abundance of DM, $\Omega_{\rm DM}h^{2}$, according to initial PBH mass, $M_{\rm in}$. The left (right) gray region is excluded from the CMB (BBN) constraint.
@@
The left panel in Fig. 2 shows the relic abundance of WIMPs due to Hawking radiation from PBHs, $\Omega_{\rm ev}h^{2}$, according to the initial PBH mass, $M_{\rm in}$, at the initial PBH density, $\beta=1\times 10^{-20}$. The left (right) gray region is excluded from the CMB (BBN) constraint. The blue solid line indicates the prediction without the memory burden effect, and the dotted lines show the predictions with the memory burden effect at q=0.0050.9q=0.005-0.9. $\Omega_{\rm FO}h^{2}\gg\Omega_{\rm ev}h^{2}$ ($\Omega_{\rm ev}/\Omega_{\rm FO}\lesssim 10^{-4}$) holds. As the relic density of WIMPs due to Hawking radiation should vanish at $t_{\rm ev}^{k}<t_{\rm FO}$, each curve has a lower limit of $M_{\rm in}$. These lower limits can indicate shifts to a lighter mass owing to the memory burden effect via $t_{\rm ev}^{k}\simeq 1/\Gamma_{\rm BH}^{k}\propto(qM_{\rm in})^{3+2k}$ from Eqs. (12) and (13).
The right panel in Fig. 2 shows the total relic abundance of DM, $\Omega_{\rm DM}h^{2}$, according to the initial PBH mass, $M_{\rm in}$. The blue solid line indicates the prediction without the memory burden effect, and the black horizontal dotted line indicates the observed relic abundance, $\Omega_{\rm DM}h^{2}=0.12$. The other dotted lines show the predictions with the memory burden effect at various $q$. The relic density of WIMPs due to FO is fixed at $\Omega_{\rm FO}h^{2}=0.08$, and $\Omega_{\rm DM}h^{2}\simeq\Omega_{\rm FO}h^{2}+\Omega_{\rm PBH}h^{2}$ holds. Hence, the trends of the dotted-line curves are reflected in the relic density of the survived PBHs with the memory burden effect, $\Omega_{\rm PBH}h^{2}\propto q\beta/M_{\rm in}^{1/2}$. The DM relic abundance is highly sensitive to the memory burden effect. For example, a small memory burden effect of $q\lesssim\mathcal{O}(10^{-3})$ is only allowed at $M_{\rm in}\lesssim\mathcal{O}(10^{6})$ g, $\beta=1\times 10^{-20}$, and $\Omega_{\rm FO}h^{2}=0.08$.
Figure 3 shows the constraint on the initial PBH density, $\beta$. The top panel shows $\beta$ against $M_{\rm in}$ at various $q$. The left (right) gray region is excluded from the CMB (BBN) constraint. The upper limit of $M_{\rm in}$ in each curve is due to the required $\beta<\beta_{\rm c}\propto 1/(q^{5/2+k}M_{\rm in}^{1+k})$.
@@
The bottom-left panel shows $\beta$ against $m_{\chi}$. As a strong memory burden effect yields long-lived PBHs, the initial PBH density, $\beta$, decreases with an increase in $q$. The upper limits of $m_{\chi}$ in the curves are obtained at $\Omega_{\rm DM}h^{2}=\Omega_{\rm FO}h^{2}$. As we fix $\Omega_{\rm DM}h^{2}=0.12$ and $\alpha_{\rm WIMP}=0.0323$, the upper limit, $m_{\chi}=858$ GeV, is obtained using Eq. (38). The lower limits of $m_{\chi}$ in the curves are controlled by the required $\beta<\beta_{\rm c}$. We obtain
$\beta\simeq\frac{\sqrt{M_{\rm in}}}{4.2\times 10^{26}q\sqrt{M_{\rm P}}}\left(\Omega_{\rm DM}h^{2}-1.7\times 10^{-10}\frac{m_{\chi}^{2}}{\alpha_{\rm WIMP}^{2}}\right),$ (41)
at $\Omega_{\rm FO}h^{2}\gg\Omega_{\rm ev}h^{2}$ from Eq.(39). Thus, a small $m_{\chi}$ with a large $q$ increases $\beta$ and eliminates it via $\beta<\beta_{\rm c}$. In principle, the WDM constraint may also contribute to the lower limit of $m_{\chi}$. However, the WDM constraint is weaker than the $\beta<\beta_{\rm c}$ constraint.
The bottom-right panel shows $\beta$ according to the coupling of WIMPs, $\alpha_{\rm WIMP}$. The lower limit of $\alpha_{\rm WIMP}$ is caused by Eq. (38). The upper limits of $\alpha_{\rm WIMP}$ in the curves are obtained from the $\beta<\beta_{\rm c}$ constraint. Because $\beta\propto m_{\chi}^{2}/\alpha_{\rm WIMP}^{2}$, the bottom-left ($\beta$ vs. $m_{\chi}$) and bottom-right ($\beta$ vs. $\alpha_{\rm WIMP}$) panels are essentially reversed.
Figure 3: Constraints on initial PBH density, $\beta$, considering WIMPs and PBHs. Top: $\beta$ vs. $M_{\rm in}$ at various $q$. The left (right) gray region is excluded from the CMB (BBN) constraint. Bottom left: $\beta$ v.s, $m_{\chi}$. Bottom right: $\beta$ vs. $\alpha_{\rm WIMP}$.
Figure 4: Relic abundance of DM and constraint on initial PBH density for FIMPs and PBHs. Only figures that differ from the case considering WIMPs are shown. Top: relic abundance of FIMPs owing to Hawking radiation from PBHs, $\Omega_{\rm ev}h^{2}$, according to initial PBH mass, $M_{\rm in}$. The left (right) gray region is excluded from the CMB (BBN) constraint. Bottom left: $\beta$ according to FIMP mass, $m_{\chi}$. Bottom right: $\beta$ according to FIMP coupling, $\alpha_{\rm FIMP}$.
@@ 4.2 FIMPs and PBHs
Next, we consider DM consisting of FIMPs and PBHs. We assume that the masses of thermal bath particles are negligible compared with those of FIMPs and that the FI mechanism is described by pair production with a constant matrix element, $|\mathcal{M}(f_{1}f_{2}\rightarrow\chi\chi)|^{2}=\alpha_{\rm FIMP}^{2}$. For a simple model of FIMPs, the relic abundance of FIMPs, $\chi$, due to FI is [111, 48]
@@
If the FI mechanism alone produces DM, then the observed relic abundance of DM, $\Omega_{\rm DM}h^{2}=0.12$, is consistent with $\alpha_{\rm FIMP}=3.52\times 10^{-11}$ [111].
According to Ref. [50], in the case of the FI mechanism, particle interactions are small enough to avoid thermalization between particles from PBHs (FIMPs) and SM particles. If PBHs dominate the universe’s energy density before evaporating, then the annihilation cross section should be several orders of magnitude larger than that in the standard FI case to ensure efficient thermalization.
In this study, PBHs never dominate the energy density in the history of the universe. DM particles from PBH evaporation are expected not to interact with thermally produced FIMPs, as in the WIMP case. To ensure that thermalization does not occur in the FI case, we compare the interaction rates between particles from PBHs and thermally produced particles with the Hubble scale [50]. However, instead of this estimation, we consider the range of parameters $q$ at $\Omega_{\rm FI}h^{2}\gg\Omega_{\rm ev}h^{2}$.
Given the feeble interactions of FIMPs, the FIMPs emitted by Hawking radiation can always contribute to the relic abundance of DM. The present-day DM density is
@@
$=9.7\times 10^{19}\alpha_{\rm FIMP}^{2}+4.2\times 10^{26}q\beta\left(\frac{M_{\rm P}}{M_{\rm in}}\right)^{1/2}$
$+2.3\times 10^{5}C\beta\frac{m_{\chi}}{\rm GeV},$ (43)
where the first–third terms are due to FI, survived PBHs, and PBH evaporation.
As in the WIMP case, we consider $\Omega_{\rm FI}h^{2}\gg\Omega_{\rm ev}h^{2}$. We regard $\alpha_{\rm FIIMP}=2.872\times 10^{-11}$, $m_{\chi}=700$ GeV ($\Omega_{\rm FI}h^{2}=0.08$), and $\beta=1\times 10^{-20}$ as the benchmark. For this benchmark, we obtain the same prediction of $\Omega_{\rm DM}h^{2}$ with $M_{\rm in}$ (right panel of Fig. 2). Thus, we omit the plot of $\Omega_{\rm DM}h^{2}$ - $M_{\rm in}$ for the FIMP case.
The top panel in Fig. 4 shows the relic abundance of FIMPs due to Hawking radiation from PBHs, $\Omega_{\rm ev}h^{2}$, according to the initial PBH mass, $M_{\rm in}$, at $\beta=1\times 10^{-20}$ and q=0.0050.9q=0.005-0.9. $\Omega_{\rm FI}h^{2}\gg\Omega_{\rm ev}h^{2}$ ($\Omega_{\rm ev}/\Omega_{\rm FI}\lesssim 10^{-4}$) holds. The left (right) gray region is excluded from the CMB (BBN) constraint.
The two bottom panels in Fig. 4 show the constraints on the initial PBH density, $\beta$, considering FIMPs. As in the WIMP case, a strong memory burden effect yields long-lived PBHs, and the initial PBH density, $\beta$, decreases with an increase in $q$ in the two bottom panels of Fig. 4.
The bottom-left panel shows $\beta$ according to the FIMP mass, $m_{\chi}$. Unlike in the WIMP case, the lower limits of $m_{\chi}$ in the curves are controlled by the WDM constraint (these lower limits are controlled by the $\beta<\beta_{\rm c}$ constraint in the WIMP case). From Eq.(43), we obtain
$\beta\simeq\frac{\sqrt{M_{\rm in}}}{4.2\times 10^{26}q\sqrt{M_{\rm P}}}\left(\Omega_{\rm DM}h^{2}-9.7\times 10^{19}\alpha_{\rm FIMP}^{2}\right),$ (44)
at $\Omega_{\rm FI}h^{2}\gg\Omega_{\rm ev}h^{2}$ in the FIMP case. Therefore, $\beta$ is independent of $m_{\chi}$ and the relation $\beta<\beta_{\rm c}$ does not constrain $m_{\chi}$. We obtain $21.2\leq m_{\chi}^{\rm min}[{\rm GeV}]\leq 40.7$ at $0.9\geq q\geq 0$ with $M_{\rm in}=1\times 10^{6}$ from the WDM constraint.
The bottom-right panel shows $\beta$ according to FIMP coupling, $\alpha_{\rm FIMP}$. Unlike in the WIMP case, the upper limit of $\alpha_{\rm FIMP}$ is caused by Eq. (42). The lower limit of $\alpha_{\rm FIMP}$ is obtained through the $\beta<\beta_{\rm c}$ condition. A small $\alpha_{\rm FIMP}$ with a large $q$ increases $\beta$ and eliminates it via the $\beta<\beta_{\rm c}$ condition.